When rats give birth, the weight of each baby varies inversely with the number of young. When there are 12 young, each baby weighs 6 grams. If a rat gives birth to 24 young, how much will each baby weigh?

Answer by stanbon(75887)

When rats give birth, the weight of each baby varies inversely with the number of young. When there are 12 young, each baby weighs 6 grams. If a rat gives birth to 24 young, how much will each baby weigh?
-----
b = k/n
Solve for "k"
6 = k/12
k = 72
-------------
Equation
wt of baby = 72/(# of babies)
Your Problem::
wt = 72/24 = 3 grams
-------------------------
Cheers,
Stan H.
-------------